OpenStudy (anonymous):

What are the intercepts of the line 5x+y=20 ? A. x- intercept:4 ; y^1 intercept: 30 B. x- intercept: ; y^1 intercept: -4 C. x- intercept: ; y^1 intercept: -20 D. x- intercept: -20 ; y-intercept: -4

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

@YoungStudier

OpenStudy (wwb00):

first convert into y=mx+b form

OpenStudy (wwb00):

to do this, you must subtract the y from both sides, then subtract the 20 from both sides.

OpenStudy (anonymous):

from what sides?

OpenStudy (wwb00):

\[5x+y=20\] \[5x+y-20=0\]

OpenStudy (wwb00):

\[5x-20=y\] \[y=5x-20\]

OpenStudy (wwb00):

substitute 0 for x to find the x intercept, then substitute 0 for y to find the y intercept.

OpenStudy (anonymous):

mm how do i find x?

OpenStudy (wwb00):

so if you put 0 in for x, \[y=5(0)-20\]

OpenStudy (wwb00):

that means the x intercept will be at -20. \[y=-20\]

OpenStudy (wwb00):

to find the y intercept, repeat the same process for y, but with a bit more work.

OpenStudy (wwb00):

\[0=5x-20\]

OpenStudy (wwb00):

\[20=5x\]

OpenStudy (wwb00):

dividing by 5, \[4=x\]

OpenStudy (wwb00):

the y intercept is at x=4

OpenStudy (wwb00):

so sorry!

OpenStudy (wwb00):

I gave you the wrong equation for the line in mx+b (standard equation)

OpenStudy (anonymous):

woops sorry restroom

OpenStudy (wwb00):

ok bear with me. all you have to do is to subsitiute 0 for x to find the x intercept, and 0 for y to find the y intercept.

OpenStudy (wwb00):

forget finding the standard form. this is the easiest way. \[5(0)+y=20\]

OpenStudy (wwb00):

\[y=20\]

OpenStudy (wwb00):

that is the x intercept.

OpenStudy (anonymous):

you mean -20?

OpenStudy (wwb00):

what are the x intercepts for b and c? I earlier stated incorrectly both the x and y intercepts. \[5x+1(0)=20\]

OpenStudy (wwb00):

substiuting 0 for y. \[5x=20\] dividing by 5 \[x=4\] I'm not getting the proper solutions.

OpenStudy (anonymous):

ok.

OpenStudy (wwb00):

wait one second.

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

back

OpenStudy (wwb00):

neither of the listed solutions are correct.

OpenStudy (wwb00):

the intercepts are at y=20 and x=4.

OpenStudy (wwb00):

I'll show you the graph.

OpenStudy (anonymous):

lets get rid of the ones that are wrong starting from A.ok?

OpenStudy (wwb00):

ok

OpenStudy (wwb00):

the y intercept can not be at 30. It must be at 20 as when x =0, y=20, thus A must be wrong.

OpenStudy (anonymous):

okay... b is next

OpenStudy (wwb00):

B. the y intercept can not be -4, as it is 20. Please look at the file I sent. C. The y intercept can not be -20, as it is positive 20. D. The y intercept can not be at -4, as it is at 20. Also, I do not see intercepts for x on B and C. Is this of the question?

OpenStudy (anonymous):

yeah

OpenStudy (wwb00):

so they are meant not to have an intercept?

OpenStudy (wwb00):

like if it were a horizontal line.

OpenStudy (anonymous):

wait on b x is 20 and on c x is -4

OpenStudy (anonymous):

my bad lol i left that out

OpenStudy (anonymous):

does that help?lol

OpenStudy (anonymous):

hello?

OpenStudy (wwb00):

so sorry.

OpenStudy (wwb00):

busy for a sec.

OpenStudy (anonymous):

itz kewl

OpenStudy (wwb00):

are you sure the intercept for A is y=30?

OpenStudy (anonymous):

noooooo its 20! im so sorry lol

OpenStudy (wwb00):

great. Its A then.

OpenStudy (wwb00):

Good luck on the rest of your HW.

OpenStudy (anonymous):

yay!thanks for the help sorry i put you through all dat xD

OpenStudy (wwb00):

It's fine. I'm glad I could help!